suppose you just got a shipment of nine televisions and 2 of them were defective. what is the probabality that both televisions work amd the probablity that at least one does not work
Answer by stanbon(75887)

suppose you just got a shipment of nine televisions and 2 of them were defective. what is the probabality that both televisions work amd the probablity that at least one does not work
-------------------------
Binomial Problem with n = 2 and p(defect) = 2/9 ; p(work) = 7/9
----
P(1<= x <=2) = 1 - P(x = 0) = (7/9)^2
=============
Cheers,
Stan H.
=========